If astronomers look toward any point in outer space, they see radiation that matches the emission spectrum of an object at 2.7 K, a remnant of the Big Bang.

a) What is the peak wavelength of this radiation? (lambda = ?)

b) What part of the electormagnetic spectum is it in?

Radio Region

Microwave Region

Visible Region

Ultraviolet Region

X-Rays Region

Gamma-Rays Region
Answer #1

From the Weins displacemnt law , the peak wave length is

lamda ( peak) T = b

where b is Weins constant

lamda ( peak) = b/T = 2.898 * 10 ^-3 K/2.7 K = 1.07 mm

(b)

micro wave region